Solving a linear system (two variables) using substitution.
Use substitution method to solve:
2x+y=-2
x-3y=-15
Thx!

Answer by MORUPHOSUOLALE(24)


2x+2y= -2--eq 1
x-3y= -15--eq 2
BY SUBSTITUTION METHOD-
If they ask you to solve by substitution method folow this step.
step 1-
In any of the equation isolate one unknown,i mean any unknown you like first.
We have two unknown i.e ,x and y in this problem.
LETS ISOLATE ANY UNKNOWN IN ANY OF THE EQUATION.
SAY equation 2 becomes-
X= -15+3y by substittion.
that means you had isolated x or in other word,you had make X equals to y and the constant.
Step 2----plug in -15+3y for x in eq(1)
2(-15+3y}+y=-2
-30+6y+y= -2
7y = -2+30
7Y= 28
Y=28/7
Y=4
To determine the value of x
plug in y=4 in any of the equation-
say eq (1)-
2X+Y= -2
2X+4= -2
2X=-2-4
2X=-6
X=-6/2
X=-3
'WE CAN ALWAYS CHECK IF OUR VALUE OF X AND Y IS CORRECT BY PLUGGING THEM IN ANY OF THE EQUATION,IF WE DO THE RIGHT HAND SIDE MUST EQUAL TO THE LEFT HAND SIDE OF THE EQUATION TO IN ORDER TO CERTIFY THAT WE ARE RIGHT.
LETS SEE-
2X+Y= -2

2(-3)+4= -2
-6+4=-2
-2=-2
This will be true for any of the equation you might have selected.
THANKS.
MORUPH OSUOLALE
CENTRAL PIEDMONT COLLEGE,
CHARLOTTE,NC